The magnitude of Coriolis accelaration for a particle of mass m moving with a velocity 100m/s horizontally along eath's surface at the north pole is about?
1
Expert's answer
2020-06-08T10:30:20-0400

We have given mass of the particle is "m" ,velocity "v" with respect to earth's frame (Non-inertial frame) is "v=100m\/s" along horizontal direction at north pole.

Let, "\Omega" is the angular velocity of earth about axis passing through center and north pole and direction is along north pole,thus "v\bot\Omega" .


Now, we know that,

Coriolis force is given by

"\overrightarrow{F}_{coriolis}=-2m( \overrightarrow\Omega\times \overrightarrow v)\\
\implies|\overrightarrow{F}_{coriolis}|=2m|( \overrightarrow\Omega\times \overrightarrow v)|\\
\implies a_{coriolis}=2|( \overrightarrow\Omega\times \overrightarrow v)|"

In this case.

"a_{coriolis}=2|( \overrightarrow\Omega\times \overrightarrow v)|=2|\overrightarrow\Omega|\cdot| \overrightarrow v|\\
=200\Omega\\
=200\times7.2921159 × 10^{-5}=14.584\times 10^{-3}m/s^2"